NEW HIGH RESOLUTION ANALYSIS OF D_{2}CO IN THE 8-12 μm REGION BY FOURIER TRANSFORM SPECTROSCOPY

\underline{A. PERRIN}, J.-M. FLAUD, Laboratoire de Photophysique Moléculaire, CNRS, Université Paris Sud, Campus d’Orsay, Bat 210, 91405 Orsay Cedex, France; A. PREDOI-CROSS, Physics Department, University of Toronto, Toronto, Ontario, Canada M5S 1A7; B. WINNEWISSER, G. MELLAU, S. KLEE, Justus Liebig Universita̋t, , Physikalisch-Chemisches Institut Heinrich-Buff-Ring 58, D-35392 Giessen, Deutschland.

A new analysis of the ν_{3}, ν_{6}, and ν_{4} interacting bands of D_{2}CO has been carried out in the 8-12 μm region using high resolution Fourier transform spectra recorded at Giessen. As compared to the previous study of these bands,^{a} improvements were obtained in the present analysis. In order to get the upper state parameters (band centers, rotational and coupling constants), the resulting ν_{3}, ν_{6}, and ν_{4} experimental energy levels were introduced in a least squares fit calculation together with the microwave measurements available in the literature. In this calculation, which allowed to reproduce both the infrared and microwave measurements within their experimental accuracies, the A- B- and C- Coriolis type interactions involving ν_{4} and ν_{6}, ν_{3} and ν_{4}, and ν_{3} and ν_{6} respectively were explicitly taken into account.

aK. Nakagawa, R. H. Schwendeman, and J.W.C. Johns, J. Mol. Spectrosc. \underline{122}, 462 (1987).